OpenStudy (anonymous):

Joy is training for a triathlon. She runs a distance of 2x miles, cycles a distance of (x^2 + 7) miles, and swims a distance of (x - 2) miles. Which simplified expression is equivalent to the total distance, in miles, Joy covers? Answers: x^2 + 3x + 5 x^2 + 3x + 9 3x^2 + 9 3x^2 + 5

OpenStudy (anonymous):

@hoblos plz help

OpenStudy (anonymous):

@romera is back

OpenStudy (hoblos):

2x + x^2 +7 +x -2 = x^2 + 3x + 5

OpenStudy (anonymous):

Do you use wolfram or something? hoblos?

OpenStudy (anonymous):

romero u are phony

OpenStudy (hoblos):

@Romero no i dont

OpenStudy (anonymous):

@hoblos he playing fake and good job in that question

OpenStudy (anonymous):

2x+x^2+7+x-2 =x^2+3x+5.....1st option

OpenStudy (anonymous):

thx @Sarkar !